A solid cylinder with a radius of 4.0 cm has the same mass as a solid sphere of radius R. If the cylinder and sphere have the same moment of inertia about their centers, what is the sphere's radius?
Answer #1
Let
Mass of the cylinder = Mass of the sphere = m
Radius of the cylinder, r = 4 cm
Radius of the sphere, R = ?
Moment of inertia of solid sphere = Moment of inertia of solid cylinder
(2/5) m R^2 = (1/2) m r^2
0.4 R^2 = 0.5 * 4^2
Radius of the sphere, R = 4.47 cm
Answer #2

The moment of inertia of solid cylinder is equal to the moment of inertia of solid sphere. 2 2 二(1.11)R, -(1.11) (4 cm) 4.47